OpenStudy (anonymous):

2.Find the quartiles for these data values: 5, 7, 7, 8, 10, 11, 12, 15, 17.

OpenStudy (anonymous):

Q1: 7 + 7 = 14 14/2 = 7

OpenStudy (anonymous):

Median: 10

OpenStudy (anonymous):

Q2: 12+15 = 27 27/2 = 13.5

OpenStudy (anonymous):

your welcome ;)